### v1

# AFABench: A Generic Framework for Benchmarking Active Feature Acquisition

Abstract.
In many real-world scenarios, acquiring all features of a data instance can be expensive or impractical due to monetary cost, latency, or privacy concerns. Active Feature Acquisition (AFA) addresses this challenge by dynamically selecting a subset of informative features for each data instance, trading predictive performance against acquisition cost. While numerous methods have been proposed for AFA, ranging from greedy information-theoretic strategies to non-myopic reinforcement learning approaches, fair and systematic evaluation of these methods has been hindered by the lack of standardized benchmarks. In this paper, we introduce AFABench, the first benchmark framework for AFA. Our benchmark includes a diverse set of synthetic and real-world datasets, supports a wide range of acquisition policies, and provides a modular design that enables easy integration of new methods and tasks. We implement and evaluate representative algorithms from all major categories, including static, greedy, and reinforcement learning-based approaches. To test the lookahead capabilities of AFA policies, we introduce a novel synthetic dataset, AFAContext, designed to expose the limitations of greedy selection. Our results highlight key trade-offs between different AFA strategies and provide actionable insights for future research. The benchmark code is available at: https://github.com/Linusaronsson/AFA-Benchmark.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
In many real-world applications, acquiring feature values of data instances can be costly. In healthcare, different medical tests can have monetary cost, latency, or privacy concerns. In recommender systems, querying user preferences can impose a cognitive burden or intrude on privacy. In such cases, acquiring the full set of features for every data instance may be infeasible or inefficient. Active Feature Acquisition (AFA) addresses this challenge by learning to dynamically select, for each data instance, a small subset of informative features to acquire, optimizing predictive performance under a constrained acquisition budget.
Unlike static feature selection, which selects the same set of features for every example, AFA enables instance-wise decisions: the features selected for one instance can differ from those selected for another. This flexibility makes AFA more realistic and efficient in domains where the informativeness or relevance of features varies between instances. Moreover, since acquisition is sequential, AFA policies can condition future acquisition decisions on previously observed feature values, mirroring expert decision-making strategies such as in clinical diagnosis, where test results guide further examination.
Despite its relevance and importance, AFA has received significantly less attention than adjacent topics such as active learning (Settles, 2009). In active learning, all features are typically available, and the goal is to selectively acquire costly labels. In contrast, AFA assumes that both labels and features are available during training, but only partial features can be acquired at test time. AFA is also different from static feature selection methods, which fix a global subset of features regardless of the test instance. For comprehensive reviews of traditional feature selection methods, we refer the reader to (Guyon and Elisseeff, 2003; Li et al., 2018; Cai et al., 2018).
A recent survey by Rahbar et al. (2025) categorizes AFA methods into several paradigms. In general, the literature can be grouped into (i) greedy methods, which acquire features one by one based on expected utility (e.g., conditional mutual information), and (ii) non-greedy methods, typically framed via reinforcement learning, which learn acquisition policies that maximize long-term rewards rather than immediate gain.
Among greedy approaches, a common theme is to select features that maximize the expected information gain. Generative methods such as EDDI (Ma et al., 2019) estimate this using partial variational autoencoders that model the joint distribution over features and targets. In contrast, discriminative methods directly estimate expected improvements in prediction loss (e.g., (Covert et al., 2023; Gadgil et al., 2024)), avoiding the need to model the entire feature distribution. Non-greedy methods model the AFA problem as a Markov Decision Process (MDP), allowing feature acquisition to be treated as sequential decision-making under budget constraints. Model-free reinforcement learning methods such as JAFA (Shim et al., 2018) directly optimize acquisition policies, while model-based methods such as ODIN (Zannone et al., 2019) leverage generative models to simulate feature acquisition trajectories. These methods can learn lookahead strategies that outperform greedy selection in certain regimes.
Although this body of work has grown considerably, progress is hindered by the lack of a standardized and unified evaluation framework. Most prior methods are evaluated in isolation using inconsistent datasets, model architectures, or acquisition costs. To our knowledge, no prior work has provided a dedicated benchmark for the AFA setting. This makes it difficult to perform fair comparisons or gain generalizable insights.
This paper introduces the first benchmark for Active Feature Acquisition. Our framework provides a unified and extensible platform for comparing AFA methods under controlled and fair conditions. It supports a diverse set of synthetic and real-world datasets and includes representative algorithms from all major categories: greedy (generative and discriminative), non-greedy (model-free and model-based RL), and static baselines. More importantly, our framework is designed to be modular and easy to extend, allowing researchers to add new methods, acquisition strategies, and datasets with minimal engineering overhead. This makes our framework a valuable tool for both researchers developing new AFA techniques and practitioners seeking to apply AFA in domain-specific tasks.
Our main contributions are:
  • We introduce AFABench, the first benchmark for Active Feature Acquisition, allowing standardized and fair comparisons across a wide range of methods and settings.
  • Our framework is modular and extensible, making it easy to add new AFA methods, datasets, and evaluation protocols.
  • We implement and evaluate representative methods from all major paradigms, including greedy (generative and discriminative), reinforcement learning-based (model-free and model-based), and static feature selection.
  • We propose a novel synthetic dataset, AFAContext, that highlights the limitations of greedy strategies and provides a testbed for evaluating non-myopic acquisition policies.
To promote reproducibility and future development, our benchmark is open source and available at: https://github.com/Linusaronsson/AFA-Benchmark. It contains step-by-step instructions on how to extend the benchmark with new models and new datasets.
2. Problem Formulation
In this section, we introduce the AFA problem and the relevant notation used throughout the paper.
2.1. Notation
Let $\mathbf{x}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{d}\}\in\mathbb{R}^{d}$ denote a data instance with $d$ features, and let $\mathbf{y}$ be the corresponding response variable in a supervised learning setting. We assume that $(\mathbf{x},\mathbf{y})$ is distributed according to a joint data distribution $p(\mathbf{x},\mathbf{y})$. Throughout, we use bold symbols (e.g., $\mathbf{x}$, $\mathbf{y}$) to represent random variables, and their non-bold counterparts (e.g., $x$, $y$) to denote specific realizations sampled from the distribution, i.e., $x,y\sim p(\mathbf{x},\mathbf{y})$. For any subset $S\subseteq[d]\triangleq\{1,\ldots,d\}$, we write $\mathbf{x}_{S}=\{\mathbf{x}_{i}\mid i\in S\}$ to denote the corresponding subset of features.
2.2. Active Feature Acquisition
As discussed in Section 1, the AFA problem appears in many variants. In this section, we describe a common formulation adopted for our benchmark.
In many machine learning settings, acquiring feature $i\in[d]$ incurs a monetary, latency, or privacy cost. AFA addresses this challenge by jointly (i) maximizing predictive performance and (ii) minimizing the total cost of queried features. Unlike static feature selection, which fixes the same feature subset for every data instance, AFA allows the set of acquired features to vary from one data instance to another, yielding a more flexible and realistic approach.
We therefore learn two functions: (i) a predictor $f(x_{S})$ that can make a prediction using any observed subset $S\subseteq[d]$; and (ii) an acquisition policy $\pi(x_{S})\in[d]\setminus S$, i.e., a function that, given the currently observed features $x_{S}$, returns the index of the next feature to reveal. They are optimized jointly via
(1) $\min_{f,\pi}\;\mathbb{E}_{p(\mathbf{x},\mathbf{y})}\!\left[\ell\!\bigl{(}f(\{\mathbf{x}_{a_{t}}\}_{t=1}^{b}),\,\mathbf{y}\bigr{)}\right],$
where $a_{n}=\pi(\{\mathbf{x}_{a_{t}}\}_{t=1}^{n-1})$ for $n=1,\dots,b$ and $b$ is a hard budget limiting how many feature values we may acquire per example. Eq. (1) adopts two simplifying assumptions: (i) all features incur the same acquisition cost, and (ii) every data instance receives an identical budget $b$. Consequently, exactly $b$ features are selected for each example. Although prior work has relaxed one or both of these assumptions, we retain them to ensure a fair and unbiased benchmark. A detailed rationale for this choice is provided in Section 3.4. However, it is straightforward to extend our framework to include these cases.
Finally, we focus the scope of our benchmark on methods developed for the common offline setting, where the goal is to deliver accurate but cost-effective predictions at test time. In this setting, access to a fully observed training dataset $\mathcal{D}=\{(x_{1},y_{1}),\dots,(x_{N},y_{N})\}$ of $N$ i.i.d. samples $(x_{i},y_{i})\sim p(\mathbf{x},\mathbf{y})$. A predictor $f$ and an acquisition policy $\pi$ are learned on $\mathcal{D}$. Then, for each test instance, the policy $\pi$ selects exactly $b$ features, and the predictor $f$ produces a prediction based solely on the acquired subset. There are several different approaches to learning $f$ and $\pi$ based on $\mathcal{D}$. In Section 3.1, we describe the methods we aim to compare in this benchmark.
3. AFABench: An Active Feature Acquisition Benchmark
In this section, we present the details of our benchmark. We begin by describing the methods included in the comparison, followed by the datasets used. We then detail the evaluation protocol used to assess the performance of each method. Finally, we outline the assumptions made to ensure a fair and unbiased evaluation.
3.1. Methods
Table 1. Summary of feature selection methods included in the benchmark. The possible categories include generative greedy ($G$), discriminative greedy (DG), model-free RL (MFRL), model-based RL (MBRL), and static selection (S).
Paradigm Strategy Greedy? Representative Method(s)
AFA (Dynamic Feature Selection) Generative estimation of CMI Greedy EDDI-GG (Ma et al., 2019)
Discriminative estimation of CMI Greedy GDFS-DG (Covert et al., 2023)
DIME-DG (Gadgil et al., 2024)
Model-free RL Non-greedy JAFA-MFRL (Shim et al., 2018)
OL-MFRL (Kachuee et al., 2019)
ODIN-MFRL (Zannone et al., 2019)
Model-based RL Non-greedy ODIN-MBRL (Zannone et al., 2019)
Oracle-based Non-Greedy AACO (Valancius et al., 2024)
Static Feature Selection Global feature importance N/A PT-S (Breiman, 2001) (Permutation tests)
CAE-S (Balin et al., 2019)
As described in the recent survey by Rahbar et al., AFA methods can be broadly categorized into two main types: (i) greedy approaches, which iteratively select the feature that provides the highest expected information gain given the features already acquired. These greedy methods can be further classified as discriminative or generative; and (ii) non-greedy approaches, which aim to learn feature selection policies typically using reinforcement learning (RL) techniques. Non-greedy methods can be further divided into model-free RL and model-based RL approaches.
In this benchmark, we have selected at least one representative state-of-the-art method from each category to provide a comprehensive evaluation of the various strategies for AFA. Table 3.1 summarizes the methods included in the benchmark and their corresponding categories. In addition, we include two static feature selection methods to highlight the potential benefits of dynamic selection. It should be noted that some existing methods do not fit neatly into any of the above categories (Rahbar et al., 2025). We discuss these methods and the rationale for excluding them from this benchmark in Section 3.4. In the following sections, we briefly describe each of the included methods.
3.1.1. Greedy methods
To quantify the information gain of a candidate feature conditioned on those already selected (as done in greedy methods), conditional mutual information (CMI) is commonly used. CMI measures the mutual dependence between a feature and the target variable, given the features observed so far. Formally, the optimal selection policy is one that chooses the feature $\mathbf{x}_{i}$ that maximizes the mutual information with the target $\mathbf{y}$, conditioned on the current set of observed features $x_{S}$. This quantity, denoted as $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$, is defined using the Kullback–Leibler divergence (Cover and Thomas, 2012; Covert et al., 2023):
(2) $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})=D_{\text{KL}}\big{(}p(\mathbf{x}_{i},\mathbf{y}\mid x_{S})\,\big{\|}\,p(\mathbf{x}_{i}\mid x_{S})\,p(\mathbf{y}\mid x_{S})\big{)}.$
The corresponding feature selection policy is then given by:
(3) $\pi(x_{S})=\operatorname*{argmax}_{i\in[n]}I(\mathbf{y};\mathbf{x}_{i}\mid x_{S}).$
In practice, implementing this ideal policy is challenging. At each decision point, computing $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$ for all feature indices $i$ requires access to the joint and conditional distributions of the features and the target, conditioned on the observed subset. In the following, we describe how the included greedy methods approximate or sidestep this computation.
EDDI-GG (Ma et al., 2019). EDDI estimates the CMI acquisition policy using a generative approach. At its core lies a novel partial Variational Autoencoder (PVAE) capable of handling arbitrarily missing features. Once trained on the full dataset $\mathcal{D}$, the PVAE offers fast approximations of posterior predictive distributions such as $p(\mathbf{x}_{i}\mid x_{S})$: it samples a latent code conditioned on the observed subset $x_{S}$ and subsequently generates the candidate feature $\mathbf{x}_{i}$. Whenever these predictive distributions admit a Gaussian parameterization, the CMI can be computed analytically; otherwise, it is efficiently estimated via Monte Carlo sampling. A limitation of this approach is that the PVAE is expensive to train. This limitation is addressed in the next two discriminative approaches.
GDFS-DG (Covert et al., 2023). This approach formulates an optimization problem that evaluates a policy by its one-step-ahead prediction accuracy. Specifically, when the policy has chosen the next feature $i$ given the currently observed set $x_{S}$, the goal is to minimize the expected loss $\mathbb{E}\!\left[\,l\bigl{(}f(x_{S},\mathbf{x}_{i}),\mathbf{y}\bigr{)}\right]$, where $l$ measures the discrepancy between the prediction and the target $\mathbf{y}$. The optimal predictor for this objective is the Bayes classifier, $f^{*}(x_{S})=p(\mathbf{y}\mid x_{S})$. Moreover, under this classifier, the policy that minimizes the expected one-step-ahead loss coincides with the CMI policy of Eq. (3). The resulting optimization is solved with amortized optimization (Amos, 2022): the variational objective is expressed in a closed form and then minimized using a deep neural network.
DIME-DG (Gadgil et al., 2024). This approach directly estimates the CMI in a discriminative way. Extending the variational estimators of Chattopadhyay et al.; Covert et al., introduces a new objective whose optimum equals the true CMI. The practical implementation uses two jointly trained networks: a predictor $f_{\theta}$ and a value network $g_{\phi}$ that outputs $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$ for each unobserved feature. When $f_{\theta}$ approaches the Bayes classifier, the expected reduction in cross-entropy achieved by adding the feature $i$ matches its CMI; training $g_{\phi}$ to reproduce this reduction yields a consistent estimator. After training, $g_{\phi}$ selects the next feature by choosing the largest predicted CMI.
3.1.2. Non-greedy methods via reinforcement learning
Given the sequential nature of active feature acquisition, framing it as a Markov Decision Process (MDP) offers a principled foundation. In this setting, reinforcement learning can be used to train non-greedy policies $\pi$ that go beyond immediate information gain, selecting features that maximize long-term reward. This approach enables the discovery of acquisition strategies that can outperform myopic and greedy methods by considering the cumulative benefit of future selections. We begin by describing the standard MDP formulation of the AFA problem for the hard-budget case (Janisch et al., 2020).
(4) $s$ $=s(S,x_{S})$
$a$ $\in U,\quad U=\{1,\dots,n\}\setminus S$
$r(s,a)$ $=r(f(x_{S}),f(x_{S\cup a}),y,\mathbb{I}(|S|=b))$
Here, $S$ are the indices of the features observed currently and $x_{S}$ are the features observed. The state $s$ is generally a function of these two vectors. In some cases, it is a simple concatenation of the two vectors, but more sophisticated methods often use special neural network architectures for encoding sets.
The discrete action space $U$ comprises the set of feature indices not yet selected. Note that we exclude the stop action often present in other papers since we focus solely on the hard-budget case.
The reward received after choosing action $a$ in state $s$, $r(s,a)$, is generally a function of the current classifier predictions, the next classifier predictions, the true label and whether the episode is finished or not. In many cases, the negative classification loss $l(f(x_{S\cup a},y)$ is used.
If one uses the reward $r(s,a)=-l(f(x_{S\cup a},y)$ without discounting, then the resulting MDP corresponds directly to the optimization objective in Eq. (1), under two simplifying assumptions: (i) a fixed budget $b$ on the number of features and (ii) uniform acquisition cost across all features (see Section 2).
JAFA-MFRL (Shim et al., 2018). This method uses the RNN-based and order-invariant set encoding introduced in Vinyals et al. to represent the subsets of acquired features, significantly reducing the complexity of the search space. The approach jointly trains a reinforcement learning agent and a classifier, ensuring that the feature selection policy is optimized with respect to the classification objective. To solve the underlying MDP, the method uses deep Q-Networks (DQN) (Mnih et al., 2015), enabling efficient policy learning even in high-dimensional feature spaces.
Adopted for the hard-budget case and with uniform acquisition costs, the reward function for this method is the negative classification loss, but only applied at the end of the episode:
(5) $r(s,a)=-l\big{(}f(x_{S\cup a}),y\big{)}\,\mathbb{I}(|S\cup a|=b).$
This sparse reward does not incentivize the agent to select features in any specific order.
In particular, this is an model-free RL approach, in contrast to the next method.
ODIN-MBRL (Zannone et al., 2019). This is a model-based RL framework for AFA. ODIN uses a pre-trained PVAE (Ma et al., 2019) to approximate the conditional distribution of unobserved features, $p(\mathbf{x}_{i}\mid x_{S})$, for $i\notin S$. Since MDP transition dynamics can be derived from this distribution, ODIN performs model-based rollouts to simulate feature acquisition trajectories, greatly improving data efficiency and robustness to missing features. The policy is trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017).
In contrast to JAFA-MFRL, this method uses the dense reward function
(6) $r(s,a)=-l\big{(}f(x_{S\cup a}),y\big{)}$
which motivates the agent to select the most informative features first.
ODIN-MFRL The key strength of ODIN-MBRL is its performance on smaller datasets (Zannone et al., 2019). However, since most of the datasets in this benchmark are relatively large, we decided to also evaluate a variant of ODIN where the agent is trained directly on the dataset, without using generative rollouts.
OL-MFRL (Kachuee et al., 2019). This is another model-free and DQN-based method. The main characteristic of it is the reward function
(7) $r(s,a)=\|\textit{Cert}(x_{S\cup a})-\textit{Cert}(x_{S})\|$
which does not depend on the true label but only on the classifier predictions. Here Cert is a vector of class probabilities, calculated by averaging over multiple forward passes with dropout.
Furthermore, the method uses a coupled PQ-network where the output of the predictor (P-network) is fed into the action-value estimator (Q-network).
3.1.3. Oracle-based methods
Rather than learning policies through reinforcement learning, oracle-based methods directly approximate an optimal acquisition policy by estimating the underlying data distributions. This approach offers a middle ground between greedy methods and computationally expensive RL training.
AACO (Valancius et al., 2024). The Acquisition Conditioned Oracle (ACO) method formulates AFA as a subset optimization problem. Instead of greedily selecting features one by one, ACO considers acquiring entire subsets of features simultaneously. The optimal policy selects the subset $o^{\prime}\subseteq\mathcal{O}$ that minimizes the expected acquisition cost plus the prediction loss,
(8) $u(x_{o},o)=\operatorname*{argmin}_{o^{\prime}\subseteq\mathcal{O}}\mathbb{E}_{y,x_{o^{\prime}}\mid x_{o}}\left[\ell\left(\hat{y}(x_{o},x_{o^{\prime}}),y\right)\right]+\alpha|o^{\prime}|,$
where $x_{o}$ denotes the currently observed features, $\mathcal{O}$ is the set of unobserved feature indices, $\hat{y}(x_{o},x_{o^{\prime}})$ is the prediction based on observed and newly acquired features, and $\alpha$ controls the acquisition cost per feature.
In practice, computing this expectation requires knowledge of the joint distribution $p(y,x_{o^{\prime}}\mid x_{o})$, which is unknown. The Acquisition Conditioned Approximate Oracle (AACO) addresses this by using $k$-nearest neighbors density estimation to approximate the required distributions. Furthermore, for computational tractability in high-dimensional spaces, AACO samples a random subset of possible feature combinations rather than evaluating all $2^{|\mathcal{O}|}$ possibilities.
This approach is non-greedy since it jointly optimizes over multiple features rather than selecting them sequentially, yet it avoids the training complexity of RL methods by directly approximating the optimal oracle policy.
3.1.4. Static feature selection methods
Static feature selection methods choose the same subset of features for all test instances, regardless of individual characteristics. We include them as baselines to highlight the potential advantage of AFA methods, which dynamically select features tailored to each instance.
PT-S (Breiman, 2001). This method involves randomly permuting the values of a single feature in all data instances and measuring the resulting drop in model performance. By disrupting the relationship between the feature and the target, this approach reveals how much the model depends on that feature. When performing permutation tests, we evaluate validation accuracy by replacing the values in each feature column with random samples drawn from the corresponding column in the training set.
CAE-S (Balin et al., 2019). This method is an end-to-end differentiable approach for global feature selection, designed to efficiently identify a subset of the most informative features while jointly training a neural network to reconstruct the input from the selected subset. Originally proposed for unsupervised feature selection via input reconstruction, we adapt it to supervised learning by modifying the prediction target, following (Covert et al., 2023).
3.2. Datasets
Table 2. Summary of datasets used in the benchmark.
Dataset Type Train Size Val Size Test Size # Features # Classes
CUBE Synthetic 700 150 150 20 8
AFAContext Synthetic 700 150 150 30 8
MNIST Real World 42,000 9,000 9,000 784 10
FashionMNIST Real World 42,000 9,000 9,000 784 10
Diabetes Real World 64,443 13,810 13,810 45 3
PhysioNet Real World 8,400 1,800 1,800 41 2
MiniBooNE Real World 91,044 19,509 19,511 50 2
In this section, we introduce the datasets used in our experiments. A summary of all datasets is provided in Table 2. CUBE is a widely used synthetic dataset in the AFA literature. We describe its structure and limitations in Section 4, and introduce a new synthetic dataset, AFAContext, to address an important shortcoming of CUBE. MNIST is the standard handwritten digit classification dataset (LeCun et al., 1998). Following prior work in AFA (Gadgil et al., 2024), we treat MNIST as a tabular task by considering each pixel as a separate feature. FashionMNIST is a dataset of grayscale images of clothing items such as shirts, trousers, and shoes (Xiao et al., 2017). It shares the same format and image dimensions as MNIST, and we similarly treat each pixel as a separate feature for tabular modeling. Diabetes is a real-world dataset for diabetes diagnosis, making it naturally suitable for AFA since different patients may require different examinations for an accurate diagnosis. This dataset has been widely used in previous AFA research (Kachuee et al., 2019; Covert et al., 2023). As in previous studies, we define three classes—normal, pre-diabetes, and diabetes—based on standard fasting glucose thresholds. The data originates from the National Health and Nutrition Examination Survey (NHANES) (NHA, 2018), a large-scale, ongoing health survey of U.S. adults and children. We use the pre-processed version made available by Kachuee et al. (2019), which includes data collected between 1999 and 2016. PhysioNet is another medical dataset, derived from the PhysioNet Challenge 2012 (Goldberger et al., 2000). It contains electronic health records (EHR) from ICU patients, with the goal of predicting in-hospital mortality, a binary classification task—based on results from various clinical tests and measurements. MiniBooNE is a particle identification dataset collected by the MiniBooNE experiment at Fermilab (Roe, 2005), where the task is to distinguish signal events (electron neutrinos) from background events (muon neutrinos) based on reconstructed particle interaction features. The dataset is fully tabular, each instance is represented by a set of real-valued detector-level features.
3.3. Evaluation Protocol
We report the performance of each method ($y$-axis) as a function of the budget, i.e., number of selected features ($x$-axis). Accuracy is used as the evaluation metric for all datasets, except PhysioNet, where we report F1-score due to its high class imbalance.
3.4. Setup for Fair Benchmarking
In this section, we discuss the assumptions necessary to provide a benchmark that is as fair and objective as possible.
Tabular data. All experiments are conducted on tabular datasets (see Section 3.2 for details), as some of the included methods do not scale well to other domains, such as image data. For example, EDDI-GG is known to not scale well beyond tabular datasets (Gadgil et al., 2024).
Hard budget. We assume a fixed feature budget $b$ for each test instance. Although many of the included methods (see Section 3.1) can be extended to the variable-budget setting—by allowing the policy $\pi$ to return a stop action indicating when to stop acquisition and make a prediction, such extensions complicate a fair comparison. For example, RL-based methods can incorporate a stop action in the action space (Janisch et al., 2020), while greedy methods might stop based on confidence thresholds, such as when the entropy of $p(\mathbf{y}\mid x_{S})$ is sufficiently low. However, these methods require tuning additional and often unintuitive hyperparameters to balance feature cost and predictive performance. Because each method handles this trade-off differently, it becomes difficult to ensure a fair comparison. Setting a fixed hard budget $b$ for all methods allows for a more controlled and objective evaluation. Moreover, specifying a fixed budget is often more practical in real-world scenarios, where users may find it easier to define a maximum number of features rather than fine-tuning cost-performance trade-offs (Janisch et al., 2020).
Common components. All methods are implemented in a unified framework, where common components between the methods are consistent. We now provide some examples of this. (i) Both EDDI-GG and ODIN-MBRL use a pre-trained PVAE (although for different purposes). We use the same pre-trained model for both. (ii) We use a shared pre-trained predictor $f$ for each task. However, some methods are specifically designed to train a predictor $f$ jointly with the policy $\pi$. In such cases, we report the performance of both the shared classifier and the one that was jointly trained.
Consistent RL framework Agents in RL-based methods interact with the same MDP, apart from the reward function, which is method-dependent.
Excluded methods. As discussed in Rahbar et al., there are other approaches to AFA that go beyond those included in our benchmark. These include (i) embedded methods and (ii) methods based on adaptive submodular objectives, such as EC2 (Golovin and Krause, 2011). Embedded methods incorporate feature acquisition directly into the model’s training or inference process—e.g., decision trees, random forests, or Naive Bayes—by modifying splitting criteria or loss functions to account for acquisition cost. However, such methods are tightly coupled to specific model architectures and cannot be employed with the shared predictor $f$ (a fully-connected neural network) used in our experiments. Methods based on adaptive submodularity focus on theoretical guarantees, but often rely on strong assumptions about the data distribution, model class, or cost structure. In contrast, we focus on more general methods that are model-agnostic and compatible with a wide range of domains and predictors.
Randomness. In our experiments, we report the average performance over (i) multiple data splits (i.e., training, validation, and test), and (ii) multiple runs with different random seeds for retraining the policy $\pi$, ensuring consistent and unified evaluation across all methods. We also include the standard deviation across runs, visualized as error bars.
Adherence to original architecture Although the most fair approach would be to perform extensive hyperparameter sweeps for each combination of methods/data sets/budgets, this is not realistic in practice. Therefore, we have opted to use the hyperparameters presented in the original paper / repository for each method, although some exceptions to this rule do exist. For example, when the parameters of a method have been tuned to datasets with a small number of features and we find that it underfits to (Fashion)MNIST, we have increased the complexity of the model. Some other changes have been made for consistency reasons. A list of the most important changes made compared to the original implementation can be found in Appendix A.
(a) CUBE
(b) AFAContext
Figure 1. Visualization of (a) CUBE, and (b) our synthetic dataset AFAContext.
4. AFAContext: A Novel Synthetic Dataset for Active Feature Acquisition
In this section, we introduce a novel dataset for AFA, called AFAContext. To motivate its design, we first describe the commonly used CUBE dataset (Rückstieß et al., 2013; Shim et al., 2018; Zannone et al., 2019), illustrated in Figure 0(a). CUBE consists of 20-dimensional real-valued vectors divided into 8 classes. For each class, three informative features are drawn from a Gaussian distribution $\mathcal{N}(\mu,0.3)$, with class-specific means $\mu$, while the remaining features are sampled from $\mathcal{N}(0.5,0.3)$ as noise. The informative features appear in different positions depending on the class, as shown by the colored cells in the figure (numbers correspond to the means). This dataset highlights the benefits of instance-specific (dynamic) feature selection, since different features are informative for different examples. However, it does not meaningfully differentiate between greedy and non-greedy methods: early in the selection process, methods must guess randomly among feature groups, and once a single informative feature is found, both greedy and non-greedy approaches can easily identify the rest. Thus, while CUBE supports evaluation of dynamic selection, it is less effective for evaluating the added value of lookahead or non-myopic strategies.
Figure 0(b) shows eight example instances from the AFAContext dataset. This dataset includes two one-hot encoded context features, $f_{1}$ and $f_{2}\in\{0,1\}$, which indicate which group of features is informative. When $f_{1}=1$, the ten features $f_{3},\dots,f_{12}$ are informative and follow the same structure as the first ten features of the CUBE dataset shown in Figure 0(a). The remaining ten features contain no information (random noise). In contrast, when $f_{2}=1$, the last ten features $f_{13},\dots,f_{22}$ are informative and also mirror the structure of CUBE. An optimal acquisition policy should therefore first select the context features $f_{1}$ and $f_{2}$ to identify the relevant context group, and then acquire the corresponding informative features. This dataset highlights both the value of instance-specific feature selection (as in CUBE) and the importance of non-greedy feature acquisition. This is because the context features $f_{1}$ and $f_{2}$ do not offer predictive value on their own and are therefore overlooked by greedy methods that focus solely on maximizing short-term performance.
5. Experiments
(a) AFAContext (Budget = 10)
(b) cube (Budget = 10)
(c) diabetes (Budget = 15)
(d) FashionMNIST (Budget = 30)
(e) physionet (Budget = 15)
(f) MNIST (Budget = 30)
(g) miniboone (Budget = 15)
Figure 2. Results for all methods across the seven datasets using the shared classifier. The reported results correspond to the large budget setting. See Appendix A.4 for other budgets and settings.
In this section, we present the experimental evaluation of the methods described in Section 3.1. Additional experimental results are illustrated in Appendix A.4. All implementation details, including model architectures and hyperparameters, are provided in Appendix A.
For each dataset, we consider three different budget values $b$ (small, medium and large) to illustrate the influence of this parameter. The choice of $b$ affects mainly the RL methods, as it is incorporated into the training process through the definition of the MDP. In addition, we conduct experiments using a shared classifier between all methods. This classifier is pretrained on the training set and then used by all methods during evaluation on the test set, ensuring a fair comparison. However, some methods include a built-in classifier that is trained jointly with the policy. To highlight the impact of this design choice, we also report results using the built-in classifier for the methods that support it.
In the main text, we present results for the large budget setting using the shared classifier (refer to the figures for the specific budget values used for each dataset). The results for all other settings are provided in the Appendix A.
Finally, for static methods, we identify the $b$ most informative features in advance on the training data. During evaluation, CAE-S randomly selects features from this set, one at a time, until the budget is exhausted, since it does not provide a relative ranking of feature importance. In contrast, PT-S selects the features sequentially in order of the estimated importance. This setup allows for a fair comparison with dynamic methods, which naturally select features one at a time.
5.1. Results on Synthetic Data
Figures 1(a)1(b) present results for the large budget setting for the two synthetic datasets, AFAContext and CUBE, respectively. For CUBE, we observe that the discriminative greedy methods, GDFS-DG and DIME-DG, achieve the strongest performance. In particular, they are the only dynamic methods that outperform the static baseline PT-S, highlighting that static methods can be highly competitive. However, given the structure of this dataset (see Figure 0(a)), it is reasonable to expect dynamic methods to outperform static ones under lower noise levels.
For AFAContext, the RL-based methods ODIN-MFRL and OL-MFRL significantly outperform all others, demonstrating that reinforcement learning approaches are well suited to capturing the non-greedy structure of this dataset. In contrast, all greedy methods perform poorly as expected. Some RL methods also show worse performance, probably due to convergence challenges. The oracle-based non-greedy method AACO also performs poorly for this dataset. This is likely due to the heuristic proposed in the original paper (Valancius et al., 2024), which enables one-at-a-time feature selection but reduces the method’s non-greedy behavior. Finally, the model-free variant of ODIN (ODIN-MFRL) outperforms its model-based counterpart (ODIN-MBRL). This suggests that model-based RL may not offer clear advantages for the AFA setting, especially when a large amount of data is available. However, there may still be scenarios where model-based training proves beneficial, depending on the dynamics of the environment or the nature of the task.
5.2. Results on Real-World Data
Figures 1(c)1(g) present results for all real-world datasets in the large budget setting. We observe that the discriminative greedy methods, GDFS-DG and DIME-DG, consistently perform among the best across all datasets. In particular, they significantly outperform other methods on FashionMNIST and MNIST. In contrast, the greedy generative method EDDI-GG performs noticeably worse. Although it aims to estimate the same underlying quantity (CMI), its performance is likely to be hindered by the limitations of the generative PVAE model used by this method. Static methods also perform well in some cases. For example, PT-S ranks among the top performers on diabetes, miniboone, and PhysioNet, suggesting that selecting the same features for all instances is effective for these datasets. On the other hand, for datasets such as MNIST and FashionMNIST, dynamic feature selection, where different features may be chosen for each instance, is substantially more beneficial.
The non-greedy methods perform relatively poorly across most of the real-world datasets. Moreover, as discussed in the previous subsection, many of the non-greedy methods perform poorly on AFAContext, where they should in principle do well. This indicates two things: (i) many real-world datasets do not exhibit non-greedy structure (as our synthetic dataset AFAContext does by construction), (ii) non-greedy methods are more difficult to train due to the increased complexity of the methods, which is typically required because finding good non-greedy policies is significantly harder than greedy selection.
Finally, we observe that AACO performs particularly poorly on FashionMNIST and MNIST. This is likely due to the limited scalability of the method in high-dimensional settings, as it relies on approximating an enumeration over all possible feature subsets by random sampling (Valancius et al., 2024). As the dimensionality increases, the quality of this approximation deteriorates, leading to reduced performance.
5.3. Compute Time Comparison
Table 3 shows the time it took to train and evaluate each method. The training time also includes pretraining time, for those methods that have such a stage. The large standard deviations are due to different datasets.
Table 3. Compute time (in seconds) comparison across methods. Values are shown as mean ± standard deviation over multiple datasets, splits, and budgets.
Method Train Time (sec) Eval. Time (sec)
JAFA-MFRL $13523\pm 2769$ $68\pm 59$
ODIN-M(F&B)RL $9271\pm 6921$ $\mathbf{23}\pm 10$
OL-MFRL $2196\pm 2267$ $\mathbf{19}\pm 6$
EDDI-$G$ $7333\pm 5823$ $2032\pm 3495$
DIME-$G$ $1697\pm 1936$ $\mathbf{19}\pm 7$
GDFS-DG $3826\pm 4395$ $\mathbf{18}\pm 6$
PT-S $268\pm 296$ $\mathbf{19}\pm 7$
CAE-S $494\pm 473$ $19\pm 7$
AACO $\mathbf{10}\pm 3$ $15676\pm 7$
RL methods generally need more computational resources while training, especially JAFA-MFRL which trains an RNN. However, their output is a simple neural network that is very fast at evaluation time. In contrast, methods like AACO require almost no training time at all but are orders of magnitude slower at inference time.
6. Conclusion
In this work, we study the problem of Active Feature Acquisition (AFA), which aims to balance predictive accuracy with the cost of acquiring features. To allow a fair and comprehensive evaluation of the AFA methods, we introduce AFABench, the first standardized benchmark framework for this task. AFABench includes a diverse collection of synthetic and real-world datasets, supports a wide range of acquisition policies, and provides a modular design to promote extensibility. We implemented and evaluated representative algorithms from major AFA paradigms—including static, greedy, and reinforcement learning-based approaches—and introduced the novel synthetic dataset AFAContext to assess the lookahead capabilities of different methods. Our empirical analysis highlights key trade-offs among AFA strategies: (i) Although non-greedy methods are capable of learning non-myopic policies that can outperform greedy selection, this is not guaranteed, largely due to the increased complexity and potential instability of these approaches. Moreover, as our results show, many real-world datasets do not exhibit strong non-greedy structure. In such cases, greedy methods, particularly discriminative ones, offer a strong alternative with significantly shorter training time. (ii) Dynamic, instance-wise feature selection is often advantageous, especially for datasets with high variability across samples. However, for some datasets, static selection performs comparably well. In addition to their competitive performance, static methods are typically more efficient to train. We hope that this benchmark will serve as a basis for future research in cost-sensitive learning and more effective feature acquisition strategies.
Acknowledgments
The work of Linus Aronsson and Morteza Haghir Chehreghani was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. The work of Valter Schütz and Morteza Haghir Chehreghani was partially supported by the Swedish Research Council VR (grant number 2023-04809). Finally, the computations and data handling was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
Appendix A Experimental Details
A.1. Implementation Details
The shared classifier used to produce the results in Section 5 is a multilayer perceptron with the following configuration:
  • Dropout rate: $0.1$
  • Learning rate: $1\times 10^{-3}$
  • Batch size: $128$
  • Two hidden layers, each with $128$ units and ReLU activations.
  • Masking probability sampled from $\mathcal{U}(0.75,0.99)$ for (Fashion)MNIST
  • Masking probability sampled from $\mathcal{U}(0,0.9)$ for the remaining datasets
It was trained until the validation loss on samples with minimum masking (i.e., probability $0.25$ on (Fashion)MNIST and 0 on the other datasets) started increasing. Then, the checkpoint where the model had the lowest validation loss was saved.
Below, we aim to fully disclose how our method implementations differ from the original implementations. Although some methods do provide original implementations, they are often poorly documented and not compatible with PyTorch (Paszke et al., 2019), which our framework is written in. In such cases, we have been forced to rewrite the method from scratch, verifying that the method works as expected on simple toy datasets.
In general, hyperparameters are always set to the values found in the paper or original implementation. In cases where a hyperparameter was not listed publicly, ”standard” values have been chosen. Examples of this include setting the learning rate to $0.001$ and $\tau=0.005$ when soft-updating target networks.
A.1.1. Common across all methods
Masking probabilities Several methods need a pretraining stage where the features are masked with varying probabilities. In our pretraining runs, we always use a fixed masking probability per batch of data, and the probability itself is sampled from the distribution
  • $\mathcal{U}(0.75,0.99)$ for MNIST and FashionMNIST
  • $\mathcal{U}(0,0.9)$ for remaining datasets
Masking all features is not useful, and since many methods can achieve reasonable performance on only very small percentage of the total number of features in (Fashion)MNIST, we increase the masking probability.
Consistent loss function We use weighted cross-entropy for all methods, using class probabilities from the training set.
A.1.2. Common across RL methods
TorchRL (Bou et al., 2023) was used extensively for the RL part. All agents were written using off the shelf loss modules, minimizing the risk for errors in the RL algorithms (like DQN and PPO). Moreover, a common AFA MDP environment was written, which all RL methods interacted with.
Computations were performed on the GPU as much as possible. This means that both the agent (with replay buffers) and the environment were placed on the GPU.
Fair number of pretraining epochs All of the RL methods have a pretraining stage, where a model is trained using supervised learning. Pretraining is always performed until convergence, always saving the model with the lowest validation loss on features masked with the minimum masking probability. As listed above, for (Fashion)MNIST this is 25% and for the other datasets it is 10%.
No discounting We do not use discounting (therefore, setting $\gamma=1$) since the MDP is episodic and the episodes are short ($\leq 30$ steps).
Consistent batch size, number of agents, and training epochs During the training stage, data is collected in consistent batches and with the same number of parallel agents for all RL methods. “Agents” in this case do not refer to separate networks, but simply to the fact that experience is collected in parallel rollouts. We use the same values as in (Shim et al., 2018), i.e.,
  • batch size $512$, with $128$ parallel agents, for all datasets except (Fashion)MNIST. This means that each agent receives a batch of size $4$.
  • batch size $64$ and $2$ parallel agents for (Fashion)MNIST. This modification is done to allow everything to fit on the GPU.
The number of batches that each agent receives is 10000 for all datasets, except for (Fashion)MNIST which needed more batches (30000).
A.1.3. JAFA-MFRL
Changes to pretraining/architecture
  • Original JAFA specifies the hidden layers for the reading block cells of the set encoder introduced in (Vinyals et al., 2016) as $[32,32]$, but not the hidden layers of the writing block. We choose to set this to $[32,32]$ as well.
  • We use $5$ processing steps for the RNN.
  • Original JAFA uses a constant masking probability of $0.5$, while ours varies.
  • Original JAFA uses hidden layers $[32,32]$ for the classifier. We do this for all datasets, except for (Fashion)MNIST where we found that the classifier severely underfitted. In this case, we increased the model complexity to $[128,128]$ instead.
Changes to RL training
  • Original JAFA uses 4-step Q-learning. Since TorchRL does not make this readily available, we instead use a TD($\lambda$) value estimator with $\lambda=0.75$. This value is chosen since it gives the largest weight to the 4-step return (see Appendix A.3).
Some additional insights gained when training this method:
  • This is by far the slowest of the RL methods, due to the RNN.
  • Using a (prioritized) replay buffer could improve the method.
  • We often found that performance (in terms of accuracy) degraded at the start of training. This could be ameliorated by only initiating joint training once $\epsilon$ for the agent has decreased.
A.1.4. ODIN-MBRL
Changes to pretraining/architecture
  • Original ODIN masks features with probability sampled from $\mathcal{U}(0,1)$, we have different probabilities.
  • The scaling factor for the KL loss term when training the PVAE is set to $0.1$ for all datasets, except for AFAContext where it is set to $0.01$. We tried the default value of $1$ first but found that it led to posterior collapse.
Changes to RL training
  • Original ODIN does not describe the number of parallel agents and the batch size that they use. We use consistent values across all RL methods.
  • In RL, it is common to perform multiple gradient steps on the same batch of data. We do a single pass, which is also what the original JAFA does.
  • For PPO we use an entropy bonus coefficient of $0.01$, the default value in TorchRL.
Original ODIN uses single linear layers for both the value network and the policy network of the PPO agent, receiving a latent vector from the PVAE as input. We found that this can cause underfitting in some cases, which is why we changed the networks to hidden layers with dimensionality $[32,32]$ for comparisons on the AFAContext dataset.
A.1.5. ODIN-MFRL
ODIN-MFRL was trained exactly the same way as ODIN-MBRL, expect that it disables generative rollouts.
A.1.6. OP-MFRL
In the original implementation, there is no pretraining stage for training the classifier separately. However, the reward that the agent receives is only as useful as the classifier it is based on, and we found that a majority of the training time was ”wasted” on waiting for the classifier to improve. Hence, we added a pretraining stage, improving training efficiency.
Changes to pretraining/architecture
  • We use the same architecture for (Fashion)MNIST as originally presented: $[512,512,128,64]$. For the remaining datasets, we use the architecture originally used in the diabetes dataset, $[64,32,16]$.
Changes to RL training
  • We do not use an adaptive learning rate.
  • Original OL originally includes the stop action, although the method is designed for the hard-budget case. We remove the stop action.
  • The OL repo only provides the number of training episodes for the diabetes dataset. As for all other RL methods, we pretrain until convergence and train with a consistent number of batches.
  • Original OL updates the target network in discrete steps, but we use a soft update instead, with $\tau=0.005$.
  • Original OL trains a single agent with batch size 1. We tried this, but it made the method very unstable. Consequently, we increased the number of agents and batch size to be consistent across all RL methods.
  • Original OL decays $\epsilon$ exponentially, but we decrease it linearly during the first half of training. This makes it consistent with our JAFA implementation.
  • Original OL prescribes a replay buffer of size $1000n_{\textrm{features}}$. We use this for all datasets except (Fashion)MNIST where the larger samples required a decreased buffer size of $30000$ to fit everything on the GPU.
Some additional insights gained when training this method:
  • This is by far the fastest of the RL methods, due to the simple architecture.
  • However, this simplicity comes at a cost: the method cannot differentiate between unobserved features and observed features with value 0. This is a significant drawback that makes the method perform especially poorly on datasets with a large number of zeros.
A.1.7. AACO
Although the AACO method avoids some problems that arise when using an RL method, to adhere to the fair benchmarking framework, several changes were needed:
  • The AACO method cannot choose a subset of features in a data instance; rather, it will break ties with selecting the feature that minimizes the loss, as per the original implementation.
  • The method cannot terminate early, as our benchmark requires a hard budget.
  • To adhere to the fair benchmark, the original classifiers suggested by Valancius et al. were not used. Instead, the shared classifier was used.
A.1.8. Common across Discriminative Greedy Methods
For both discriminative greedy methods, we use the same predictor structure to ensure fair comparison. Specifically, we use a predictor with two hidden layers of size [128,128] and a dropout rate of 0.3. Both methods involve predictor pretraining and joint training of the predictor and selector/estimator. During training, we concatenate the masked features with the feature mask as the model input according to the default settings of (Covert et al., 2023; Gadgil et al., 2024). We use a batch size of 128 and a learning rate of 0.001 during both model pretraining and training, and we use early stopping in both cases.
A.1.9. GDFS-DG
We use the default gradual temperature annealing strategy (Covert et al., 2023) to train the models during the joint training phase. During model training, we gradually decrease the temperature from 1.0 to 0.1, with a total of five temperature stages using logarithmic spacing. At each temperature stage, we train the model for a maximum of 250 epochs. We use the model with the lowest validation loss at each temperature as the starting point for the next temperature stage and will continue to train until the best model is obtained in the final stage.
A.1.10. DIME-DG
During the joint training phase of DIME-DG, we perform feature selection using the $\epsilon$-greedy strategy with an initial $\epsilon$ set to 0.05. If the validation accuracy does not improve over five consecutive epochs, we reduce the $\epsilon$ by multiplying it by 0.2, up to a maximum of ten times. We also use the default bounded strategy (Gadgil et al., 2024) for CMI estimation. We jointly train the model for a maximum of 250 epochs.
A.1.11. EDDI-GG
After joint pretraining of partial variational autoencoder and predictor, the EDDI-GG model does not require further training. Hence, we only save the same models for different budget values in the training stage and use this model for subsequent evaluation.
A.1.12. Common across Static Methods
During the training phase of the static method, we use a selector and classifier with two hidden layers of size [128,128] for both methods. We train both methods with a batch size of 128 and a learning rate of 0.001. We train each method for a maximum of 250 epochs with an early stopping strategy.
A.1.13. CAE-S
To ensure fair comparison,instead of selecting the optimal features for each budget value from 1 to the maximum feature budget, we use the CAE selector to only select the optimal features for the maximum budget. We then use this selected feature list to train a separate predictor for each budget.
A.2. Hardware
All experiments were performed on NVIDIA T4 GPUs.
A.3. Proof that $\lambda=0.75$ gives maximum weight to 4-step return
The $\lambda$-return that TD($\lambda$) (Sutton, 1988) estimates is
(9) $G_{t}^{\lambda}\doteq(1-\lambda)\sum_{n=1}^{\infty}\lambda^{n-1}G_{t:t+n}$
where $G_{t:t+n}$ is the n-step return,
(10) $G_{t:t+n}\doteq R_{t+1}+\gamma R_{t+2}+\dots+\gamma^{n-1}R_{t+n}+\gamma^{n}Q(S_{t+n},A_{t+n}).$
The weight given to the 4-step return $G_{t:t+4}$ in $G_{t}^{\lambda}$ is
(11) $w_{4}=(1-\lambda)\lambda^{3}$
which when differentiated,
(12) $\frac{\partial}{\partial\lambda}w_{4}=\lambda^{2}(3-4\lambda)$
shows that $w_{4}$ is maximized by $\lambda=0.75$. $\qquad\qquad\qquad\qquad\square$
A.4. Additional Results
This section supplements the experimental results in different budget settings, covering the three budget ranges. We evaluate the performance of different methods using both the builtin classifiers and the shared classifiers, where the results are illustrated in Figures˜A.1, A.2, A.3, A.4 and A.5. These experiments are mainly conducted to evaluate the robustness of our conclusions in different experimental settings.
From the experimental results, we observe that the overall trends are consistent with the findings in the main paper under the large budget setting, with the relative performance rankings of each method mostly remaining stable. We also observe that in most cases the results using builtin classifiers are slightly better than those using shared classifiers. This outcome is expected, as the builtin classifier is usually jointly optimized with the feature selection strategy.
We also observe that with a small budget, reinforcement learning methods tend to select more informative features earlier, resulting in behaviors that more closely resemble greedy methods. In contrast, for large budgets, their selection strategy is relatively more balanced. This observation confirms our initial hypothesis that the behavior of reinforcement learning methods is affected by budget constraints. As an example, the JAFA-MFRL method achieves $\approx 55\%$ accuracy with budget $10$ on FashionMNIST when evaluated with the shared classifier (see Figure 1(d)). In contrast, the same method only obtains $\approx 45\%$ accuracy after $10$ features when trained with budget $30$ on the same dataset (see Figure 1(d)).
(a) AFAContext (Feature Budget = 5)
(b) cube (Feature Budget = 5)
(c) diabetes (Feature Budget = 10)
(d) FashionMNIST (Feature Budget = 20)
(e) physionet (Feature Budget = 10)
(f) MNIST (Feature Budget = 20)
(g) miniboone (Feature Budget = 10)
Figure A.1. Performance of Different Methods on Seven Datasets Using Shared Classifiers (Medium Budgets)
(a) AFAContext (Feature Budget = 3)
(b) cube (Feature Budget = 3)
(c) diabetes (Feature Budget = 5)
(d) FashionMNIST (Feature Budget = 10)
(e) physionet (Feature Budget = 5)
(f) MNIST (Feature Budget = 10)
(g) miniboone (Feature Budget = 5)
Figure A.2. Performance of Different Methods on Seven Datasets Using Shared Classifiers (Small Budgets)
(a) AFAContext (Feature Budget = 10)
(b) cube (Feature Budget = 10)
(c) diabetes (Feature Budget = 15)
(d) FashionMNIST (Feature Budget = 30)
(e) physionet (Feature Budget = 15)
(f) MNIST (Feature Budget = 30)
(g) miniboone (Feature Budget = 15)
Figure A.3. Performance of Different Methods on Seven Datasets Using Builtin Classifiers (Large Budgets)
(a) AFAContext (Feature Budget = 5)
(b) cube (Feature Budget = 5)
(c) diabetes (Feature Budget = 10)
(d) FashionMNIST (Feature Budget = 20)
(e) physionet (Feature Budget = 10)
(f) MNIST (Feature Budget = 20)
(g) miniboone (Feature Budget = 10)
Figure A.4. Performance of Different Methods on Seven Datasets Using Builtin Classifiers (Medium Budgets)
(a) AFAContext (Feature Budget = 3)
(b) cube (Feature Budget = 3)
(c) diabetes (Feature Budget = 5)
(d) FashionMNIST (Feature Budget = 10)
(e) physionet (Feature Budget = 5)
(f) MNIST (Feature Budget = 10)
(g) miniboone (Feature Budget = 5)
Figure A.5. Performance of Different Methods on Seven Datasets Using Builtin Classifiers (Small Budgets)